# Guided Flow Policy: Learning from High-Value Actions in Offline Reinforcement Learning

### v1

@@ 4.1 Main results: extensive offline RL benchmarks
Benchmarks. We conducted extensive experiments over a suite of robot locomotion and manipulation tasks, spanning three major benchmarks: D4RL fu2020d4rl, its successor Minari minari, and the recently proposed OGBench park2024ogbench. For comparability with existing works, we first evaluate on D4RL’s AntMaze (6 tasks) and Adroit (12 tasks). We also present results on Minari, evaluating both GFP and FQL, to facilitate the community’s migration from D4RL. Minari includes all available Gym-Mujoco datasets (Hopper, HalfCheetah, and Walker, each with 3 tasks), as well as Adroit (12 tasks). Our most extensive evaluation focuses on OGBench, which offers substantially more complex and challenging tasks than D4RL. Following park2025flow, we use the reward-based single-task variants of OGBench. We evaluate across 9 locomotion and 11 manipulation environments, each with 5 tasks, for a total of 100 state-based tasks, and we also consider 5 pixel-based tasks. Combined with D4RL and Minari, we benchmark GFP on 144 tasks overall, including evaluation of prior methods, we did about 15 000 runs in total. Our JAX-based implementation of GFP will be released after the rebuttal phase. It can complete one training run in under 30 minutes on modern GPUs.
@@
Comparison against previous works. We first compare GFP against 10 prior methods, including standard offline RL, flow, and diffusion-based methods on the 50 OGBench tasks reported by park2025flow (see Sec. 5 and Fig 5 for a presentation of prior works, Tab. 10 for the full results). Performance profiles synthesizing these comparisons are presented Fig. 3(a), following agarwal2021deep. GFP clearly stands out compared to all these prior works. We also compare against 6 methods on D4RL (Tab. 11) and evaluate both GFP and FQL on Minari (see Tab. 12) to facilitate the community’s migration from D4RL.
@@
Extensive study on 144 tasks. GFP is further evaluated against 3 baselines: (i) FQL park2025flow, as it comes second only to GFP on the first 50 tasks; (ii) ReBRAC tarasov2023revisiting, as we were able to improve its performance compared to previously reported results substantially, see Sec. 4.3, and (iii) IQL kostrikov2021offline to represent in-sampling approaches. Tab. 2 summarizes our results grouped by environment, with detailed per-task results in the appendix (Tabs. 8, 9, 11, and 12). Together with performance profile plots Figs. 3(b) and 3(c), it confirms that GFP achieves state-of-the-art performance, with particularly substantial gains on noisy and challenging environments. For instance, on the cube-double-noisy and cube-triple-noisy datasets, GFP achieves an average score of 63 and 24, respectively, compared to 38 and 4 for FQL, and 20 and 5 for ReBRAC. Similarly, GFP stands out in some very challenging locomotion tasks, such as humanoidmaze-large-navigate (18 vs. 7 for FQL and 13 for ReBRAC), and manipulation tasks, such as cube-triple-play (16 vs. 4 for FQL and 3 for ReBRAC).
@@
Figure 3: OGBench analysis. (a) Performance profiles for 50 tasks comparing GFP against a wide range of prior works, showing the fraction of tasks where each algorithm achieves a score above threshold $\tau$, using the evaluation reported by park2025flow. (b) Performance profiles on 105 tasks, including more challenging ones, and carefully reevaluated prior methods. (c) Performance profiles restricted to 30 noisy and explore tasks.
@@
Table 2: Offline RL results. GFP achieves best or near-best performance on all 144 benchmark tasks. Results are averaged over 8 seeds for state-based tasks and 4 seeds for pixel-based ones, with values reported from prior works park2025flow; tarasov2023revisiting; fu2020d4rl in italic, and values within 95% of the best performance are shown in bold. GFP actor $\pi_{\theta}$ represents our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of the training procedure. Full per-task results are provided in the appendix Tabs. 8, 9, 11, and 12, and the comparison with 10 methods on the 50 previously evaluated OGBench tasks Tab. 10.
@@ 4.2 Analysis of the temperature parameter $\eta$
We investigate the impact of the temperature-controlled guidance mechanism. In Fig. 4, GFP is evaluated with varying temperature $\eta$ on two challenging noisy tasks from OGBench, characterized as highly suboptimal data according to park2024ogbench. The presence of low-quality demonstrations makes selective action emphasis decisive for effective learning. Figs. 4(a) and 4(c) demonstrate the advantages of moderate temperatures. Very low temperatures cause training instability due to over-concentration on narrow action sets, while very high temperatures fail to provide sufficient filtering of suboptimal actions. As the temperature decreases, VaBC $\pi_{\omega}$ performance improves, confirming that the value-guided filtering mechanism successfully emphasizes higher-value actions, until the temperature is too low.
@@ 4.3 Re-evaluation of prior works on OGBench
To obtain a fair comparison of GFP against prior methods, we reevaluate existing baselines on OGBench. During the development of our method, we observed that several task-specific hyperparameters (e.g., discount factor $\gamma$, minibatch size $B$, and critic aggregation scheme for doubled Q-learning fujimoto2018addressing) have a significant impact on performance. Tab. 3 reports results under these revised settings, showing that careful tuning can substantially improve the reported scores of both ReBRAC and FQL. Since the optimal values for these hyperparameters were generally consistent across methods, we treated them as task-specific and, by default, applied the same settings to GFP, FQL, and ReBRAC (see Tabs. 4 and 5 in the appendix for detailed values).
@@
Task Environment Bigger discount factor $\gamma$ Previously reported park2025flow Our evaluations ReBRAC FQL ReBRAC FQL antmaze-large-navigate (5 tasks) $\mathit{81}\pm{\scriptstyle 5}$ $\mathit{79}\pm{\scriptstyle 3}$ $95.9\pm{\scriptstyle 0.4}$ $88.1\pm{\scriptstyle 3.4}$ humanoidmaze-large-navigate (5 tasks) $\mathit{2}\pm{\scriptstyle 1}$ $\mathit{4}\pm{\scriptstyle 2}$ $12.9\pm{\scriptstyle 4.2}$ $6.5\pm{\scriptstyle 2.7}$ Bigger minibatch size Previously reported park2025flow Our evaluations ReBRAC FQL ReBRAC FQL antmaze-giant-navigate (5 tasks) $\mathit{26}\pm{\scriptstyle 8}$ $\mathit{9}\pm{\scriptstyle 6}$ $33.2\pm{\scriptstyle 5.7}$ $16.3\pm{\scriptstyle 8.2}$ Same critic aggregation for ReBRAC as used in FQL Previously reported park2025flow Our evaluations ReBRAC ReBRAC humanoidmaze-medium-navigate (5 tasks) $\mathit{22}\pm{\scriptstyle 8}$ $59.2\pm{\scriptstyle 12.1}$ antsoccer-arena-navigate (5 tasks) $\mathit{0}\pm{\scriptstyle 0}$ $55.9\pm{\scriptstyle 1.5}$ cube-double-play (5 tasks) $\mathit{12}\pm{\scriptstyle 1}$ $12.6\pm{\scriptstyle 1.8}$ scene-play (5 tasks) $\mathit{41}\pm{\scriptstyle 3}$ $41.6\pm{\scriptstyle 3.6}$ puzzle-4x4-play (5 tasks) $\mathit{14}\pm{\scriptstyle 1}$ $17.1\pm{\scriptstyle 1.3}$
@@ 5 Related Work
Offline RL. Early methods addressed distributional shift by working on both how the critic is learned and how the policy is extracted. On the critic side, Conservative Q-Learning (CQL) kumar2020conservative penalizes out-of-distribution value estimates and Implicit Q-Learning (IQL) kostrikov2021offline avoids querying such values via expectile regression, see Fig 5. For policy extraction, weighted behavior cloning treats offline RL as supervised learning, typically using the advantage value as the weight, as in AWR or AWAC peng2019advantage; nair2020awac, see Fig 5. The policy extraction method found to lead to the best performance and scalability in standard offline-RL is behavior-regularized policy gradient fujimoto2021minimalist; park2024value, abbreviated DDPG+BC and also called reparametized policy gradients, which consists in directly maximizing the value function, while regularizing with a behavior cloning term. The BRAC family studied in this paper uses behavior-regularized policy gradient and trains the critic directly using the actor, see Fig 5. ReBRAC tarasov2023revisiting demonstrated state-of-the-art performance from this simple idea, through careful hyperparameter tuning and architectural choices.
@@
Control with diffusion and flow models. Expressive generative models enable multi-modal action and trajectory distributions modeling and have been integrated into control along several axes. They have been used for trajectory and motion planning, trajectory optimization, hierarchical control  janner2022planning; ajay2022conditional; zheng2023guided; liang2023adaptdiffuser; li2023hierarchical; suh2023fighting; venkatraman2023reasoning; lee2023refining; ma2024hierarchical; chen2024simple; pan2024model; le2025sobolev, as well as for world modeling ding2024diffusionworld; alonso2024diffusion; lee2024gta. In offline RL, numerous ideas emerged to integrate diffusion and flow models in the various policy extraction methods.
@@
Weighted behavior cloning. A straightforward approach is to add weight in front of the cloning loss of diffusion and flow models, examples include FAWAC park2025flow, EDP kang2023efficient, and QVPO ding2024diffusionbased. Related energy-guided methods explicitly target to sample from $\pi_{\theta}(a\mid s)\propto\mu(a\mid s)\exp(\eta Q(s,a))$, where $\mu$ is the behaviour policy. This includes CEP lu2023contrastive and QIPO zhang2025energy; alles2025flowq, see Fig 5.
@@
Behavior-regularized policy gradient. Combining the most effective policy extraction method with expressive models is a motivating research direction: Diffusion-QL wang2022diffusion (Fig 5), DiffCPS he2023diffcps, trust-region formulations for diffusion policies (DTQL) chen2024diffusion, entropy Q-ensemble regularization zhang2024entropy, Consistency-AC with consistency models ding2023consistency, and diffusion policy gradients li2024learning; yang2023policy. In that case, the main challenge is the drawbacks of backpropagation through time (BPTT) when trying to directly optimize the value of the outputs of an iterative process. To avoid BPTT, FQL park2025flow (Fig 5) trains a flow matching with a standard BC loss, and then leverages it in the BC regularization term of a one-step BRAC actor. Beyond these families, other policy extractors include rejection sampling from a fixed BC generative policy: IDQL hansen2023idql (Fig 5), DICE mao2024diffusion, and SfBC chen2022offline.
@@
Positioning GFP within prior work. Our method, GFP (illustrated Fig 1 and Fig 5), extends the BRAC family of offline RL algorithms and is closely related to FQL park2025flow, sharing the same goal of avoiding backpropagation through time. However, the distinctive contribution of GFP lies in integrating the weighted BC approach for the regularization component of the BRAC actor. Rather than using weighted BC to train the final actor directly (as in Fig 5), GFP employs it to inject value-awareness into the regularizer VaBC. Compared to other methods using a weighted-BC policy, VaBC is not GFP’s actor, its weights are computed using the one-step actor trained by DDPG+BC, with the latter using VaBC as a regularizer. This bidirectional guidance between the two policies fundamentally distinguishes GFP from other methods that use weighted BC as their main actor, computing the weights based on the weighted BC policy itself, see Fig 5.
Removing the value-aware guidance from GFP flow policy would yield FQL (Fig 5), and removing the one-step actor would imply using the flow policy to train the critic, yielding FAWAC (Fig 5). Alternatively, one could use rejection sampling instead of weighted-BC to regularize the actor with the best actions found by a BC policy, yet at the cost of sampling multiple actions from the flow policy for each policy update step.
@@ 6 Discussion and Conclusion
In this work, we revisited behavior regularization for offline RL. BRAC approaches have been shown to be highly effective for policy extraction tarasov2023revisiting; park2024value. However, conventional BRAC approaches constrain the learned policy to remain close to the raw dataset distribution. While this reduces instability, the regularization term itself typically does not distinguish between the high and low-value actions. This limitation is problematic in suboptimal datasets, where regularizing with all transitions indiscriminately can hinder performance.
@@ Acknowledgments
This work has received support from the French government, managed by the National Research Agency, under the France 2030 program with the references Organic Robotics Program (PEPR O2R) and “PR[AI]RIE-PSAI” (ANR-23-IACL-0008). This research was funded, in part, by l’Agence Nationale de la Recherche (ANR), projects RODEO (ANR-24-CE23-5886) and PEPR O2R – PI3 ASSISTMOV (ANR-22-EXOD-0004). The European Union also supported this work through the ARTIFACT project (GA no.101165695) and the AGIMUS project (GA no.101070165). The Paris Île-de-France Région also supported this work in the frame of the DIM AI4IDF. The authors gratefully acknowledge the support and resources provided by the CLEPS infrastructure at Inria Paris. This work was performed using HPC resources from the GENCI-IDRIS Jean-Zay cluster (Grant 2024-AD010616763). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the funding agencies.

### v2

@@ 4.1 Main results: extensive offline RL benchmarks
Benchmarks. We conducted extensive experiments over a suite of robot locomotion and manipulation tasks, spanning three major benchmarks: D4RL (Fu et al., 2020), its successor Minari (Younis et al., 2024), and the recently proposed OGBench (Park et al., 2024a). For comparability with existing works, we first evaluate on D4RL’s AntMaze (6 tasks) and Adroit (12 tasks). We also present results on Minari, evaluating both GFP and FQL, to facilitate the community’s migration from D4RL. Minari includes all available Gym-Mujoco datasets (Hopper, HalfCheetah, and Walker, each with 3 tasks), as well as Adroit (12 tasks). Our most extensive evaluation focuses on OGBench, which offers substantially more complex and challenging tasks than D4RL. Following Park et al. (2025), we use the reward-based single-task variants of OGBench. We evaluate across 9 locomotion and 11 manipulation environments, each with 5 tasks, for a total of 100 state-based tasks, and we also consider 5 pixel-based tasks. Combined with D4RL and Minari, we benchmark GFP on 144 tasks overall, including evaluation of prior methods, we did about 15 000 runs in total. Our JAX-based implementation of GFP will be released after the rebuttal phase. It can complete one training run in under 30 minutes on modern GPUs.
@@
Comparison against previous works. We first compare GFP against 10 prior methods, including standard offline RL, flow, and diffusion-based methods on the 50 OGBench tasks reported by Park et al. (2025) (see Sec. 5 and Fig 5 for a presentation of prior works, Tab. 10 for the full results). Performance profiles synthesizing these comparisons are presented Fig. 3(a), following Agarwal et al. (2021). GFP clearly stands out compared to all these prior works. We also compare against 6 methods on D4RL (Tab. 11) and evaluate both GFP and FQL on Minari (see Tab. 12) to facilitate the community’s migration from D4RL.
@@
Extensive study on 144 tasks. GFP is further evaluated against 3 baselines: (i) FQL (Park et al., 2025), as it comes second only to GFP on the first 50 tasks; (ii) ReBRAC (Tarasov et al., 2023), as we were able to improve its performance compared to previously reported results substantially, see Sec. 4.3, and (iii) IQL (Kostrikov et al., 2021) to represent in-sampling approaches. Tab. 2 summarizes our results grouped by environment, with detailed per-task results in the appendix (Tabs. 8, 9, 11, and 12). Together with performance profile plots Figs. 3(b) and 3(c), it confirms that GFP achieves state-of-the-art performance, with particularly substantial gains on noisy and challenging environments. For instance, on the cube-double-noisy and cube-triple-noisy datasets, GFP achieves an average score of 63 and 24, respectively, compared to 38 and 4 for FQL, and 20 and 5 for ReBRAC. Similarly, GFP stands out in some very challenging locomotion tasks, such as humanoidmaze-large-navigate (18 vs. 7 for FQL and 13 for ReBRAC), and manipulation tasks, such as cube-triple-play (16 vs. 4 for FQL and 3 for ReBRAC).
@@
Figure 3: OGBench analysis. (a) Performance profiles for 50 tasks comparing GFP against a wide range of prior works, showing the fraction of tasks where each algorithm achieves a score above threshold $\tau$, using the evaluation reported by Park et al. (2025). (b) Performance profiles on 105 tasks, including more challenging ones, and carefully reevaluated prior methods. (c) Performance profiles restricted to 30 noisy and explore tasks.
@@
Table 2: Offline RL results. GFP achieves best or near-best performance on all 144 benchmark tasks. Results are averaged over 8 seeds for state-based tasks and 4 seeds for pixel-based ones, with values reported from prior works (Park et al., 2025; Tarasov et al., 2023; Fu et al., 2020) in italic, and values within 95% of the best performance are shown in bold. GFP actor $\pi_{\theta}$ represents our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of the training procedure. Full per-task results are provided in the appendix Tabs. 8, 9, 11, and 12, and the comparison with 10 methods on the 50 previously evaluated OGBench tasks Tab. 10.
@@ 4.2 Analysis of the temperature parameter $\eta$
We investigate the impact of the temperature-controlled guidance mechanism. In Fig. 4, GFP is evaluated with varying temperature $\eta$ on two challenging noisy tasks from OGBench, characterized as highly suboptimal data according to Park et al. (2024a). The presence of low-quality demonstrations makes selective action emphasis decisive for effective learning. Figs. 4(a) and 4(c) demonstrate the advantages of moderate temperatures. Very low temperatures cause training instability due to over-concentration on narrow action sets, while very high temperatures fail to provide sufficient filtering of suboptimal actions. As the temperature decreases, VaBC $\pi_{\omega}$ performance improves, confirming that the value-guided filtering mechanism successfully emphasizes higher-value actions, until the temperature is too low.
@@ 4.3 Re-evaluation of prior works on OGBench
To obtain a fair comparison of GFP against prior methods, we reevaluate existing baselines on OGBench. During the development of our method, we observed that several task-specific hyperparameters (e.g., discount factor $\gamma$, minibatch size $B$, and critic aggregation scheme for doubled Q-learning (Fujimoto et al., 2018)) have a significant impact on performance. Tab. 3 reports results under these revised settings, showing that careful tuning can substantially improve the reported scores of both ReBRAC and FQL. Since the optimal values for these hyperparameters were generally consistent across methods, we treated them as task-specific and, by default, applied the same settings to GFP, FQL, and ReBRAC (see Tabs. 4 and 5 in the appendix for detailed values).
@@
Task Environment Bigger discount factor $\gamma$ Previously reported in Park et al. (2025) Our evaluations ReBRAC FQL ReBRAC FQL antmaze-large-navigate (5 tasks) $\mathit{81}\pm{\scriptstyle 5}$ $\mathit{79}\pm{\scriptstyle 3}$ $95.9\pm{\scriptstyle 0.4}$ $88.1\pm{\scriptstyle 3.4}$ humanoidmaze-large-navigate (5 tasks) $\mathit{2}\pm{\scriptstyle 1}$ $\mathit{4}\pm{\scriptstyle 2}$ $12.9\pm{\scriptstyle 4.2}$ $6.5\pm{\scriptstyle 2.7}$ Bigger minibatch size Previously reported in Park et al. (2025) Our evaluations ReBRAC FQL ReBRAC FQL antmaze-giant-navigate (5 tasks) $\mathit{26}\pm{\scriptstyle 8}$ $\mathit{9}\pm{\scriptstyle 6}$ $33.2\pm{\scriptstyle 5.7}$ $16.3\pm{\scriptstyle 8.2}$ Same critic aggregation for ReBRAC as used in FQL Previously reported in Park et al. (2025) Our evaluations ReBRAC ReBRAC humanoidmaze-medium-navigate (5 tasks) $\mathit{22}\pm{\scriptstyle 8}$ $59.2\pm{\scriptstyle 12.1}$ antsoccer-arena-navigate (5 tasks) $\mathit{0}\pm{\scriptstyle 0}$ $55.9\pm{\scriptstyle 1.5}$ cube-double-play (5 tasks) $\mathit{12}\pm{\scriptstyle 1}$ $12.6\pm{\scriptstyle 1.8}$ scene-play (5 tasks) $\mathit{41}\pm{\scriptstyle 3}$ $41.6\pm{\scriptstyle 3.6}$ puzzle-4x4-play (5 tasks) $\mathit{14}\pm{\scriptstyle 1}$ $17.1\pm{\scriptstyle 1.3}$
@@ 5 Related Work
Offline RL. Early methods addressed distributional shift by working on both how the critic is learned and how the policy is extracted. On the critic side, Conservative Q-Learning (CQL) (Kumar et al., 2020) penalizes out-of-distribution value estimates and Implicit Q-Learning (IQL) (Kostrikov et al., 2021) avoids querying such values via expectile regression, see Fig 5. For policy extraction, weighted behavior cloning treats offline RL as supervised learning, typically using the advantage value as the weight, as in AWR or AWAC (Peng et al., 2019; Nair et al., 2020), see Fig 5. The policy extraction method found to lead to the best performance and scalability in standard offline-RL is behavior-regularized policy gradient (Fujimoto and Gu, 2021; Park et al., 2024b), abbreviated DDPG+BC and also called reparametized policy gradients, which consists in directly maximizing the value function, while regularizing with a behavior cloning term. The BRAC family studied in this paper uses behavior-regularized policy gradient and trains the critic directly using the actor, see Fig 5. ReBRAC (Tarasov et al., 2023) demonstrated state-of-the-art performance from this simple idea, through careful hyperparameter tuning and architectural choices.
@@
Control with diffusion and flow models. Expressive generative models enable multi-modal action and trajectory distributions modeling and have been integrated into control along several axes. They have been used for trajectory and motion planning, trajectory optimization, hierarchical control  (Janner et al., 2022; Ajay et al., 2022; Zheng et al., 2023; Liang et al., 2023; Li et al., 2023; Suh et al., 2023; Venkatraman et al., 2023; Lee et al., 2023; Ma et al., 2024; Chen et al., 2024a; Pan et al., 2024; Le Hellard et al., 2025), as well as for world modeling (Ding et al., 2024b; Alonso et al., 2024; Lee et al., 2024). In offline RL, numerous ideas emerged to integrate diffusion and flow models in the various policy extraction methods.
@@
Weighted behavior cloning. A straightforward approach is to add weight in front of the cloning loss of diffusion and flow models, examples include FAWAC (Park et al., 2025), EDP (Kang et al., 2023), and QVPO (Ding et al., 2024a). Related energy-guided methods explicitly target to sample from $\pi_{\theta}(a\mid s)\propto\mu(a\mid s)\exp(\eta Q(s,a))$, where $\mu$ is the behaviour policy. This includes CEP (Lu et al., 2023) and QIPO (Zhang et al., 2025; Alles et al., 2025), see Fig 5.
@@
Behavior-regularized policy gradient. Combining the most effective policy extraction method with expressive models is a motivating research direction: Diffusion-QL (Wang et al., 2022) (Fig 5), DiffCPS (He et al., 2023), trust-region formulations for diffusion policies (DTQL) (Chen et al., 2024b), entropy Q-ensemble regularization (Zhang et al., 2024), Consistency-AC with consistency models (Ding and Jin, 2023), and diffusion policy gradients (Li et al., 2024; Yang et al., 2023). In that case, the main challenge is the drawbacks of backpropagation through time (BPTT) when trying to directly optimize the value of the outputs of an iterative process. To avoid BPTT, FQL (Park et al., 2025) (Fig 5) trains a flow matching with a standard BC loss, and then leverages it in the BC regularization term of a one-step BRAC actor. Beyond these families, other policy extractors include rejection sampling from a fixed BC generative policy: IDQL (Hansen-Estruch et al., 2023) (Fig 5), DICE (Mao et al., 2024), and SfBC (Chen et al., 2022).
@@
Positioning GFP within prior work. Our method, GFP (illustrated Fig 1 and Fig 5), extends the BRAC family of offline RL algorithms and is closely related to FQL (Park et al., 2025), sharing the same goal of avoiding backpropagation through time. However, the distinctive contribution of GFP lies in integrating the weighted BC approach for the regularization component of the BRAC actor. Rather than using weighted BC to train the final actor directly (as in Fig 5), GFP employs it to inject value-awareness into the regularizer VaBC. Compared to other methods using a weighted-BC policy, VaBC is not GFP’s actor, its weights are computed using the one-step actor trained by DDPG+BC, with the latter using VaBC as a regularizer. This bidirectional guidance between the two policies fundamentally distinguishes GFP from other methods that use weighted BC as their main actor, computing the weights based on the weighted BC policy itself, see Fig 5.
Removing the value-aware guidance from the GFP flow policy yields FQL (Fig 5), and removing the one-step actor implies using the flow policy to train the critic, yielding FAWAC (Fig 5). Alternatively, one could use rejection sampling instead of weighted-BC to regularize the actor with the best actions found by a BC policy, yet at the cost of sampling multiple actions from the flow policy for each policy update step.
@@ 6 Discussion and Conclusion
In this work, we revisited behavior regularization for offline RL. BRAC approaches have been shown to be highly effective for policy extraction (Tarasov et al., 2023; Park et al., 2024b). However, conventional BRAC approaches constrain the learned policy to remain close to the raw dataset distribution. While this reduces instability, the regularization term itself typically does not distinguish between the high and low-value actions. This limitation is problematic in suboptimal datasets, where regularizing with all transitions indiscriminately can hinder performance.
@@ Acknowledgments
This work has received support from the French government, managed by the National Research Agency, under the France 2030 program with the references Organic Robotics Program (PEPR O2R) and “PR[AI]RIE-PSAI” (ANR-23-IACL-0008). This research was funded, in part, by l’Agence Nationale de la Recherche (ANR), projects NIMBLE project (ANR-22-CE33-0008), RODEO (ANR-24-CE23-5886), PEPR O2R - AS2 (ANR-22-EXOD-0006), and PEPR O2R – PI3 ASSISTMOV (ANR-22-EXOD-0004). The European Union also supported this work through the ARTIFACT project (GA no. 101165695) and the AGIMUS project (GA no. 101070165). The Paris Île-de-France Région also supported this work in the frame of the DIM AI4IDF. The authors gratefully acknowledge the support and resources provided by the CLEPS infrastructure at Inria Paris. This work was performed using HPC resources from the GENCI-IDRIS Jean-Zay cluster (Grant 2024-AD010616763). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the funding agencies.